OpenStudy (anonymous):

Which of the following is the solution to the equation 252c = (4c – 10) ?

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

252c - 4c = -10 248c = -10 -10 / 248 = c

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

is the same as -5/124 is approximately equal to -.040322 etc.

OpenStudy (anonymous):

Which of the following is the solution to the equation 252c = (4c – 10) ? c = – c = c = 4 c = 7

OpenStudy (anonymous):

c - 5/2 c= 7/2

OpenStudy (anonymous):

one of those

OpenStudy (anonymous):

you are putting the equation in here incorrectly or something because -5/124 is the answer

OpenStudy (anonymous):

can you wait 1 sec so i can type it right

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

25^2c= square root 5^(4c-10)

OpenStudy (anonymous):

is this better

OpenStudy (turingtest):

rewrite 25 as a power of 5 then use exponent rules to simpilify and get a common base of 5 between everything

OpenStudy (turingtest):

simplify*

OpenStudy (anonymous):

idk how

OpenStudy (anonymous):

im trying to learn for a 400 point test so i studying

OpenStudy (turingtest):

\[25=5^2\]